Reports of 19787 Clear search Modify search
DetChar (General)
shoichi.oshino - 9:14 Tuesday 09 August 2022 (21739) Print this report
SummaryPage trouble: process was stopped from 8/8 7:00 UTC
SummaryPage GIF process was stopped from 8/8 7:00 UTC.
I deleted this process.
After this work, SummaryPage starts to create plots.
DetChar (General)
shoichi.oshino - 9:11 Tuesday 09 August 2022 (21738) Print this report
SummaryPage trouble: process was stopped from 8/7 7:00 UTC
SummaryPage VIS process was stopped from 8/7 7:00 UTC.
I deleted this process.
After this work, SummaryPage starts to create plots.
CRY (Cryostat IY)
shinji.miyoki - 7:54 Tuesday 09 August 2022 (21736) Print this report
Comment to Turned (21710)

The IYV temp keeps gradually increasing.

Images attached to this comment
CAL (XPcal)
dan.chen - 6:16 Tuesday 09 August 2022 (21735) Print this report
Integration sphere position marking and optical efficiency measurement

Date: 2022/8/8

With Takashi Takase

We checked beam alignment (link), and adjusted WSK (old name: GSK) position against the Pcal beams, then marked the positions for higher reproducibility during the integration sphere calibration.

Details: link

CAL (XPcal)
dan.chen - 6:15 Tuesday 09 August 2022 (21734) Print this report
XPcal beam position check

Date: 2022/8/8

I checked XPcal beam positions on ETM, and the positions were in acceptable regions.

Details: link

MIF (General)
kenta.tanaka - 20:28 Monday 08 August 2022 (21732) Print this report
Comment to Turn on PZT drivers and add PZT offsets for POP STM1 and STM2 (21730)

## Abstract

I calculated the beam position displacement on ITM with ABCDEF matrix. According to PZT specsheet, PZT applied at 75V pushes the mirror forward 1.8 um (= 3.6um/150V*75V). When STM1 and STM2 pushed the mirror forward by 1.8um each, the beam spot position on the ITM shifted ~0.1 mm from the position before pushing.

### How to calculate it

  • The propagation of rays in a system slightly misaligned from the ABCD system is calculated with the matrix called "ABCDEF matrix", which is an extension of the original 2x2 ABCD matrix to a 3x3 matrix by adding a shift of the beam position due to misalignment to the elements (1,3), a tilt of the optical axis due to misalignment to the elements (2,3), and (0,0,1) in the third row.
  • The ABCDEF matricies of free space propagation, reflection from a displaced and misaligned mirror and thin lens is represented as below.
    • free propagation L^free=[1l0010001]\hat{L}_{\mathrm{free}} = \left[ \begin{matrix} 1 & l & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
    • reflection from a mirror R^mirror=[102Δxsinθ-2Re12Δα001]\hat{R}_{\mathrm{mirror}} = \left[ \begin{matrix} 1 & 0 & 2\Delta x \sin \theta \\ -\frac{2}{R_e} & 1 & 2\Delta \alpha \\ 0 & 0 & 1 \end{matrix} \right]
    • thin lens F^lens=[100-1f10001]\hat{F}_{\mathrm{lens}} = \left[ \begin{matrix} 1 & 0 & 0 \\ -\frac{1}{f} & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
      • where l is distance in the free space, Δx\Delta x is lateral shift of a mirror, θ\theta is an incident angle to a mirror, ReR_e  is an effective radious of curvature of a mirror, Δα\Delta \alpha is a misaligned angle of a mirror and ff is a focus length of a lens (c.f. Siegman "Lasers" chap. 15)
  •  The ABCDEF matrix from STM1 to ITM is as below
    •  M^=L^pr3itmR^pr3L^pr2pr3L^pr2GrL3F^GrL3L^GrL3GrM3R^GrM3L^GrM3GrM2R^GrM2\hat{M} = \hat{L}_{\mathrm{pr3itm}} \hat{R}_{\mathrm{pr3}} \hat{L}_{\mathrm{pr2pr3}} \hat{L}_{\mathrm{pr2GrL3}} \hat{F}_{\mathrm{GrL3}} \hat{L}_{\mathrm{GrL3GrM3}} \hat{R}_{\mathrm{GrM3}} \hat{L}_{\mathrm{GrM3GrM2}} \hat{R}_{\mathrm{GrM2}}
  • Assuming that the ray vector before shifting is r0=[00]r_0 = \begin{bmatrix} 0 \\ 0 \end{bmatrix}, and that STM1 and STM2 are shifted forward by 1.8um (= 3.6 um /150 V * 75 V) each and no misalignment has occurred, the beam shift on the ITM is equal to M(1,3).
    • M(1,3) ~ 0.1 mm
    • R_e
MIF (General)
takafumi.ushiba - 19:23 Monday 08 August 2022 (21730) Print this report
Turn on PZT drivers and add PZT offsets for POP STM1 and STM2

Related work with klog21163.

[Toriyama, Ushiba]

Abstract:
We turned on PZT drivers and added 75 V offsets for each PZTs from DGS.
Alignment seems to come back after adding all PZT offset, so we didn't change anything else.

Detail:
I implemented EUL2PZT matrices for STM1 and STM2 as shown in fig1.
Then, I implemented calibration factor from PZT driver output voltage to corresponding DGS cnts in K1:ALS-X_IAL_STM{1,2}_PZT filter banks (fig2).
Since 0V (0 cnts) - 10V (2^15 cnts) are corresponding to 0V - 150V, a calibration factor is (2^15)/150 = 218.45 cnts/V.
Thanks to this calibration factors, we can set the offset of the filter banks as the voltage that we would like to apply to the PZTs.


We connected PZT drivers to AC power supply, which is used for PDA1 of fiber output monitors and prepared inside the POP table.
After that, we turned on PZT drivers and set analog outputs from the driver to zeros.


First, we added offsets for PZTs of STM2 one by one.
Figure 3 shows the time series data of transmitted GRX signals during the work.
As you can see, transmitted light power once go down but come back after all offsets were added.

In the same way, we aded offsets for PZTs of STM1 one by one.
Figure 4 shows the time trend of  transmitted GRX signals during the work.
In this time, we cannot see any significant drop of the signals.

After that, I requested the XARM guardian to ALS_LOCKED state and then we can lock XARM with TEM00 (fig5).
Since the transmitted light power didn't change so much (~1.5: fig6), we didn't perform any additional alignment workafter adding PZT offsets.

Note:
As I wrote, PZT drivers are connected to the same power supply for PDA1 of fiber output monitor now.
I'm not sure it shold be avoided but even if so, we don't have to hesitate changing PZT driver power supply because we confirmed PZTs don't affect the alignment so much if we move all of them at the same time.

Tanaka-kun kindly calculated how much beam spot change is expected due to this work and it seems negligible.
Detail will be reported by Tanaka-kun.

Images attached to this report
Comments to this report:
kenta.tanaka - 20:28 Monday 08 August 2022 (21732) Print this report

## Abstract

I calculated the beam position displacement on ITM with ABCDEF matrix. According to PZT specsheet, PZT applied at 75V pushes the mirror forward 1.8 um (= 3.6um/150V*75V). When STM1 and STM2 pushed the mirror forward by 1.8um each, the beam spot position on the ITM shifted ~0.1 mm from the position before pushing.

### How to calculate it

  • The propagation of rays in a system slightly misaligned from the ABCD system is calculated with the matrix called "ABCDEF matrix", which is an extension of the original 2x2 ABCD matrix to a 3x3 matrix by adding a shift of the beam position due to misalignment to the elements (1,3), a tilt of the optical axis due to misalignment to the elements (2,3), and (0,0,1) in the third row.
  • The ABCDEF matricies of free space propagation, reflection from a displaced and misaligned mirror and thin lens is represented as below.
    • free propagation L^free=[1l0010001]\hat{L}_{\mathrm{free}} = \left[ \begin{matrix} 1 & l & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
    • reflection from a mirror R^mirror=[102Δxsinθ-2Re12Δα001]\hat{R}_{\mathrm{mirror}} = \left[ \begin{matrix} 1 & 0 & 2\Delta x \sin \theta \\ -\frac{2}{R_e} & 1 & 2\Delta \alpha \\ 0 & 0 & 1 \end{matrix} \right]
    • thin lens F^lens=[100-1f10001]\hat{F}_{\mathrm{lens}} = \left[ \begin{matrix} 1 & 0 & 0 \\ -\frac{1}{f} & 1 & 0 \\ 0 & 0 & 1 \end{matrix} \right]
      • where l is distance in the free space, Δx\Delta x is lateral shift of a mirror, θ\theta is an incident angle to a mirror, ReR_e  is an effective radious of curvature of a mirror, Δα\Delta \alpha is a misaligned angle of a mirror and ff is a focus length of a lens (c.f. Siegman "Lasers" chap. 15)
  •  The ABCDEF matrix from STM1 to ITM is as below
    •  M^=L^pr3itmR^pr3L^pr2pr3L^pr2GrL3F^GrL3L^GrL3GrM3R^GrM3L^GrM3GrM2R^GrM2\hat{M} = \hat{L}_{\mathrm{pr3itm}} \hat{R}_{\mathrm{pr3}} \hat{L}_{\mathrm{pr2pr3}} \hat{L}_{\mathrm{pr2GrL3}} \hat{F}_{\mathrm{GrL3}} \hat{L}_{\mathrm{GrL3GrM3}} \hat{R}_{\mathrm{GrM3}} \hat{L}_{\mathrm{GrM3GrM2}} \hat{R}_{\mathrm{GrM2}}
  • Assuming that the ray vector before shifting is r0=[00]r_0 = \begin{bmatrix} 0 \\ 0 \end{bmatrix}, and that STM1 and STM2 are shifted forward by 1.8um (= 3.6 um /150 V * 75 V) each and no misalignment has occurred, the beam shift on the ITM is equal to M(1,3).
    • M(1,3) ~ 0.1 mm
    • R_e
MIF (ITF Control)
masayuki.nakano - 14:56 Monday 08 August 2022 (21729) Print this report
Comment to ALS DARM was locked (21659)

OMG!! I missed a really great news!!!

Congrats!!!
I was really surprised by your speed to succeed in the type-A control, and I'm pretty sure that you will lock FPMI and be ready for the noise hunting!!

Go KAGRA, go!!

FCL (Air)
shinji.miyoki - 11:13 Monday 08 August 2022 (21728) Print this report
BS-C cooler stopped arpund 9:30

Hayakawa-kun found BS-C cooler stopped around 9:30(It worked around 8:00). The reason could be the reduction of water flow because of the cryo-cooler operation for IYC duct shields on Friday.

So, we should make a set of checks when we operated something that requires water in the corner station because the continuous operation of the Daikin coolers in the corner station is so sensitive to the amount of water flow.

CRY (Cryostat IY)
shinji.miyoki - 10:58 Monday 08 August 2022 (21727) Print this report
Comment to Turned (21710)

I turned off 2 of 10 FFUs in the IYC clean booth. 8 are now operated.

VIS (IY)
ryutaro.takahashi - 10:20 Monday 08 August 2022 (21726) Print this report
Offload of BF GAS

The BF GAS COILOUTF has been saturated due to the cooling. I offloaded the BF GAS with the FR by +21000step.

CRY (Cryostat IY)
takafumi.ushiba - 9:02 Monday 08 August 2022 (21725) Print this report
Comment to Turned (21710)

[Shiota, Ushiba]

Before starting cryocoolers, we centered MN RPY and V OpLevs.

CRY (Cryostat IY)
shinji.miyoki - 7:27 Monday 08 August 2022 (21724) Print this report
Comment to Turned (21710)

The IYV/C temp seemed to be up by 0.1C.

Images attached to this comment
VIS (General)
takafumi.ushiba - 16:11 Sunday 07 August 2022 (21723) Print this report
High-bandwidth angular control of Type-A OpLev DC control

Abstract:
I changed OpLev DC control for all Type-A suspensions.
In the new control, UGFs for pitch and yaw are set as 1.6 Hz and 0.4 Hz, respectively.
New controls are implemented to the guardian, so users don't need to take care of the changes.

Detail:
I measured transfer functions from MN_TEST_{P,Y} to MN_OLDAMP_{P,Y}_IN1 for all Type-A suspension.
Measured data was store at /users/ushiba/measurement/TF/TypeA/20220807/

After that, I redesigned OpLev DC control to increase the UGF of DC controls.
In the old design, UGFs are set as 0.2 and 0.08 Hz for pitch and yaw, respectively.
In the new design, these are changed to 1.6 and 0.4 Hz, respectively.

Thanks to this modification, micro seismic band (~0.2 Hz) is inside the conrol band.
So, it can supress 0.2 Hz angualr oscillation during ALS DARM lock due to microseismic motions.

Note:
During this work, I noticed dewhitening filters for TM OpLev signals of ITMY and ETMY was engaged, which should be not engaged.
So, I disabled them.

VIS (SR3)
takafumi.ushiba - 16:01 Sunday 07 August 2022 (21722) Print this report
Hierarchical angular control for SR2

I implemented hierarchical anglar control for SR3 in the same manner as BS's (klog21662).

VIS (SR2)
takafumi.ushiba - 16:00 Sunday 07 August 2022 (21721) Print this report
Hierarchical angular control for SR2

I implemented hierarchical anglar control for SR2 in the same manner as BS's (klog21662).

VIS (IMM)
takafumi.ushiba - 16:23 Saturday 06 August 2022 (21720) Print this report
Comment to High bandwidth OpLev control for IMMT suspensuions (21718)

I increased UGF of yaw OpLev DC control of IMMTs from 1 Hz to 3 Hz because there is a peak of yaw resonance in the OpLev signals.

DGS (General)
takahiro.yamamoto - 14:13 Saturday 06 August 2022 (21719) Print this report
maintenance day in a while
It's a maintenance day for the first time in three weeks.

Security updates were applied for k1gate and gwdet.
If some troubles will be found, please let me know.

Some updates of real-time models was done for adding DQ channels.
- k1asc was updated for adding POP forward related signals as the DQ channel which had been prepared in klog#21566.
- For the IFO calibration by Michelson fringes, some RFPD signals are added as DQ channels with 256Hz sampling on k1lsc.
- Duration for capturing TCam movie can be changed from a MEDM screen. It's defined on k1grdconfig.

By the way, for clearing SDF Diffs, roughly 1 hour is needed. It's more than 10% of the maintenance time...
It takes a great deal of time to make a decision of accepting or reverting, especially if it is not written in the klog and the corresponding person of these changes is not in the control room to hear about it soon.
Let's manage SDF frequently even if a maintenance day is skipped.
VIS (IMM)
takafumi.ushiba - 20:18 Friday 05 August 2022 (21718) Print this report
High bandwidth OpLev control for IMMT suspensuions

Abstract:
I implemented high-bandwidth (~ 1Hz) angular control for IMMTs.
It seems working welll.

Detail:
I measured mechanical transfer functions of IMMTs and made suspension models.
Measured data was stored at /kagra/Dropbox/Measurement/VIS/PLANT/{IMMT1,IMMT2}/2022/08/.
Reference number of the measurement is 202208051908.
Suspension models are stored at FM10 (SUSMod_220805) of TM_OLDAMP filter banks.

After making the model, I designed new OpLev DC control filter.
In this time, UGF is set at 1 Hz, which is 100 times larger than the previous design.
I tested new DC control and it seems working well.

Note:
OpLev DC controls are still disengaged at LOCK_ACQUISITION state because we need to change the gain of ASC control if we use setpoint pass for ASC.

Comments to this report:
takafumi.ushiba - 16:23 Saturday 06 August 2022 (21720) Print this report

I increased UGF of yaw OpLev DC control of IMMTs from 1 Hz to 3 Hz because there is a peak of yaw resonance in the OpLev signals.

VIS (IMM)
takafumi.ushiba - 20:09 Friday 05 August 2022 (21717) Print this report
Comment to IMMT1 pico channel was different from MEDM discription (21678)

I checked picomotor movement (only ch1) and confirmed pitch can be moved by using ch1 now.

VAC (EY)
nobuhiro.kimura - 17:02 Friday 05 August 2022 (21716) Print this report
Leak hunting for GVey
When: 02/Aug./2022
Who: Kimura, Takahashi (ICRR) and Yasui
Task:
We attempted to identify and repair a leak in the gate valve enclosure on August 2.
As a result, we identified the leak as connection #5 in the figure. We injected "Vacseal" diluted to 5% with acetone into the leak,
but this did not stop the leak.
Thus, a leak of 1x10^9 Pam^3/s still remains.
The location of the leak and the commercial potential of the leak reaction are shown in the attached drawings and photographs.
Images attached to this report
VIS (General)
satoru.ikeda - 16:14 Friday 05 August 2022 (21715) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

DGS Regular maintenance day(8/5)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO
  Check if the script can be started
  Checked that the limit switches installed are green.
 [Check result]
  No problem
  ITMX is suspended for maintenance at other locations.

Release Sensor
 [Checked the Release Sensor.]
 Checked that Release Sensor is 0, 1: Seated.
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Confirmation]
  PingTest is performed.
 [Result]
  Only known

 NG: 
  SR3(IM)
   => Hang. Rebooted.
  
  MCI,STM1,
  AS_WFS,POP_POM,POS_POM,POM1
  MCF(MCO,MCI)
   => Confirmation only, no recovery operation was performed.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(IM,BF,STM)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,IMMT2,OMMT1,OMMT2,OSTM
  REFL_WFS,POP,POP2,POS,POS2
  PCAL(EX1,EX2,EY1,EY2):Not checked because it is ON only when used

Temperature controller
 [Checked]
 Is the temperature acquisition working properly?

 [Check result] 
  Is the temperature acquisition working properly?
  => No problem

Precision air Processer
 [Check Contents]
 Is the temperature acquisition working properly?

 [Result]
  No problem

FunctionGenerator
 (Confirmation)
 Is there a ping response?
 Are all EPICS channels alive?
  [sitemap]-[Commissioning]-[ALS FG]-[Xarm],[Yarm].
 [Check result]
 X and Y PLLs are working properly.

HWP control PC
 (Confirmation)
 Check if it is possible to login to the control PC.
 (Result) 
 Both k1hwp0 and 1 are OK.
 

Images attached to this comment
Non-image files attached to this comment
VIS (IMM)
koji.nakagaki - 15:49 Friday 05 August 2022 (21714) Print this report
Comment to IMMT1 pico channel was different from MEDM discription (21678)
[Ikeda, Nakagaki]

We checked the wiring of the driver at the site and found that the cables for ch1 and ch2 were plugged in incorrectly.
(cable 2 to port 1, cable 1 to port 2)
We corrected them to the correct connections.

There was a handwritten label on the cable, but it had almost disappeared, so the label was reattached.
Images attached to this comment
DGS (General)
satoru.ikeda - 15:26 Friday 05 August 2022 (21713) Print this report
Comment to LAN cable replacement (21690)

LAN Maintenance in the Control Room

Background

 Many LAN packet errors are occurring at k1ctr1, k1ctr2, and k1ctr4.
 Upon investigation, we found that the Z620 LAN used in the WS has two ports, but the errors occur when connected to the upper LAN port.

 Upon investigation, both were Intel NICs, but the interfaces connected seem to be different.
 Error tolerance under high load appears to be different.

Correspondence.

 For additional investigation, we connected the LAN cable connected to the upper LAN port to the lower LAN port (AMT LAN).
 We also modified /kagra/bin/SYS/netUSE.sh and restarted deamon.

 Targets are the following WS.
  k1ctr1, k1ctr2, k1ctr4
  k1ctr25, which uses Z620, was also modified.

Addition

 In conjunction with this, a defective LAN cable was replaced.

 The target is the following WS.
 LAN cables of k1mon,k1epics(Fig1) were replaced.

Other investigations showed that k1naoj02 and k1naoj07 also had many packet errors.
There is no fix proposed here.
 

Images attached to this comment
Non-image files attached to this comment
MIF (ASC)
hirotaka.yuzurihara - 14:39 Friday 05 August 2022 (21712) Print this report
Comment to Script to reset ASC feedback is available (21675)

Based on Ushiba-san's request, I added the `ALL` option in Script to reset ASC feedback is available (The script is same one). I added new button on medm screen of ALLOPLEVS. (see the attached photo)

By pushing the ALL button in `offload oplev 2`, the script will reset ASC feedback for ITMX, ETMX, ITMY, ETMY, PR2, PR3, IMMT1, IMMT2, SR3 one by one. The current ramp time is set as 3 seconds. To finish ALL, it will take ~27seconds.

Images attached to this comment
Search Help
×

Warning

×